OpenStudy (anonymous):

Jack tries to define a function g: ℚ → ℚ by the rule g(m/n) = m - n for all integers m and n with n ≠ 0. Susan claims g is not well defined. Who is right? Justify your answer

OpenStudy (anonymous):

what is \(g(\frac{8}{12})\)?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

nooo -4

OpenStudy (anonymous):

what is \(g(\frac{4}{6})\)?

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

and finally what is \(g(\frac{2}{3})\)?

OpenStudy (anonymous):

okkk so its not well defined

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

clear right?

OpenStudy (anonymous):

yupppp i got that. Thanks that was awesome grandpa

OpenStudy (anonymous):

i prefer "old head" now i think i will put on an episode of perry mason or perhaps put on a van morrison album

OpenStudy (anonymous):

hahahahaah